OpenStudy (anonymous):

find the derivative of F(x)=integral from x to 4 of (sin(t^(4)))dt

OpenStudy (anonymous):

\[\int_x^4sin(t^4)dt?\]

OpenStudy (anonymous):

well what is \[\frac{d}{dx}\int\]

OpenStudy (anonymous):

yeah that^

OpenStudy (anonymous):

use the fundamental theorem of calculus

OpenStudy (anonymous):

\[f(x)=\frac{d}{dx}F(x)\]

OpenStudy (anonymous):

\[F(x)=\int_x^4sin(t^4)dt\]

OpenStudy (anonymous):

Okay, gimme a sec.

OpenStudy (anonymous):

I find F(x) very vague for this i think it should be \[F(x)=\int_x^4 sin(t^4)=F(4)-F(x) \] \[\frac{d}{dx}(F(4)-F(x))=0-F'(x)(1)=-f(x)\]

OpenStudy (anonymous):

the one inside is F(t) for x whereas the outside is just F(x)

OpenStudy (anonymous):

so it ends up being \[f(t)=sin(t^4)\] \[f(x)=f(t)|_0^x\] you can say i guess

OpenStudy (anonymous):

\[-sin(x^4)\]

OpenStudy (anonymous):

make sense?